OpenStudy (anonymous):

Find the slope of a line that is a) parallel and b) perpendicular to the given line: -4x-9y=-4

OpenStudy (anonymous):

a) 4/9 b)-9/4

OpenStudy (anonymous):

y=4/9-4x/9

OpenStudy (anonymous):

How did you figure it out?

OpenStudy (anonymous):

you plug in the values.

OpenStudy (saifoo.khan):

Solve for y: Then compare it with y=mx + b where "m" is the slope.

OpenStudy (anonymous):

4x+ 9y=4 -> 9y = -4x + 4 -> y = -4x/9 + 4/9 slope m = -4/9 Slope of parallel line = - 4/9 Slope of perpendicular line = +9/4